What is the formula of that compound in which 50% x element and 50% y element of total mass?
1
Expert's answer
2017-07-05T10:55:06-0400
Let the total mass is 100 g.
n(X)=(50 g)/(10 g/mole)=5 mol.
n(Y)=(50 g)/(20 g/mole)=2.5 mol.
Simplest atomic ratio:
n(X)/n(Y) =5/2.5=2.
The formula of that compound is
X_2 Y.